Solar radiation from the sun (T=5500K) heats an asphalt driveway on a summer day. The driveway has dimensions of 4.0 x 15 m. If the asphalt driveway behaves like a black body with emissivity of 1,

a)determine the amount of heat that the driveway absorbs in 1 hour.?

b) If the surface temperature of the driveway in part (a) reaches a temperature of 38C, determine how much heat it will emit in 1 hour. Is your answer the same or different as in part a)?


1
Expert's answer
2021-08-02T08:41:10-0400

a)

"A_1=Pt=\epsilon St=(1+\kappa)\sigma T_1^4 abt=22.4\cdot 10^{12}~J,"

b)

"A_2=(1+\kappa)\sigma T_2^4 abt=2.3\cdot 10^8~J."